OpenStudy (anonymous):

how many moles are in 9.45*10^15 molecules of o3?

OpenStudy (anonymous):

Ok, so this is O3 as in ozone? A mole of a compound/substance(I dunno the right word), by definition, is measured as in 6.02x10^23 of it's molecule/atom. So let's calculate: \[9.45\times 10^{23} \times \frac{ 1 mole }{ 6.02 \times 10^{23} } = ?\]